I really don't understand ratios (I really need help with this problem). Twelve of the 27 students in the class are boys. What is the ratio of girls to boys in the class? can you put scratchwork with it to please and thankyou
Answer by stanbon(75887)

I really don't understand ratios (I really need help with this problem). Twelve of the 27 students in the class are boys. What is the ratio of girls to boys in the class?
-----------------------------
# of girls = 27-12 = 15
---
Ratio of girls to boys = 15/12 = 5/4 or 5:4
===========
Cheers,
Stan H.
============